# Sparse Probabilistic Synthesis of Quantum Operations

### v1

@@ IV Discussion and conclusion
The present work is an important stepping stone in developing advanced quantum protocols whereby the quantum resources are significantly reduced through the use of classical computers that solve non-trivial but offline, classical pre and post processing tasks. While related probabilistic techniques are routinely used in the context of quantum error mitigation, the present work significantly generalises these mathematical concepts to a wide spectrum of entirely new application scenarios in quantum technologies.
@@ Acknowledgments
B.K. thanks the University of Oxford for a Glasstone Research Fellowship and Lady Margaret Hall, Oxford for a Research Fellowship. The numerical modelling involved in this study made use of the Quantum Exact Simulation Toolkit (QuEST), and the recent development QuESTlink [39] which permits the user to use Mathematica as the integrated front end, and pyQuEST [40] which allows access to QuEST from Python. We are grateful to those who have contributed to all of these valuable tools. The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility [41] in carrying out this work and specifically the facilities made available from the EPSRC QCS Hub grant (agreement No. EP/T001062/1). The author also acknowledges funding from the EPSRC projects Robust and Reliable Quantum Computing (RoaRQ, EP/W032635/1) and Software Enabling Early Quantum Advantage (SEEQA, EP/Y004655/1).
@@ A.1 Quantum error mitigation
For this reason we consider applications where one assumes perfect unitary operations and the only source of error is due to, e.g., finite T depth or discretisation of quantum rotation angles etc. Furthermore, operations $\mathcal{R}_{desired}$ and $\mathcal{U}_{l}$ are known exactly, hence arbitrary precision numerical representations can be used. This is all in contrast to QME [27] whereby one approximately learns these from experiments and one thus needs to consider low-precision numerical representations.
@@ A.2 Probabilistic gate synthesis
Probabilistic gate sequences have been employed to improve the accuracy of Clifford+T sequences [42, 43, 44]. In particular, the desired unitary operation is approximated as $\mathcal{U}_{desired}\approx\sum_{l}p_{l}\mathcal{U}_{l}$ where $p_{l}$ is a probability distribution and $\mathcal{U}_{l}$ are is a set of operations close to $\mathcal{U}_{desired}$ but with finite T depth. It was observed that the approximation error in $\sum_{l}p_{l}\mathcal{U}_{l}$ is always quadratically smaller than in $\mathcal{U}_{l}$. However, these techniques are approximate and as one increases the number of gates in the circuit, the fidelity of the quantum state decreases exponentially as demonstrated in [24].
@@ B.1 Proof of 1
$\mathbb{E}(\hat{\mathcal{R}}_{desired})=\sum_{l}p(l)\hat{\mathcal{R}}_{desired}=\sum_{l}p(l)\lVert\gamma\rVert_{1}\text{sign}(\gamma_{l})\mathcal{U}_{l},$
@@
$\mathbb{E}(\hat{\mathcal{R}}_{desired})=\sum_{l}\gamma_{l}\text{sign}(\gamma_{l})\mathcal{U}_{l}=\mathcal{R}_{desired}$
@@ B.3 Unbiasedness
B.4 Bounding the number of shots

### v2

@@ IV Discussion and conclusion
The present work is a promising new direction in developing advanced quantum protocols whereby quantum resources are significantly reduced through the use of classical computers that solve non-trivial but offline, classical pre- and post-processing tasks. While related probabilistic techniques are routinely used in the context of quantum error mitigation, the present work extends these mathematical concepts to a spectrum of new application scenarios in quantum technologies.
Data availability: An example code is available online at https://github.com/BalintKoczor/probsynth.
@@ Acknowledgments
BK thanks UKRI for the Future Leaders Fellowship project titled Theory to Enable Practical Quantum Advantage (MR/Y015843/1). BK thanks the University of Oxford for a Glasstone Research Fellowship and Lady Margaret Hall, Oxford for a Research Fellowship. The numerical modelling involved in this study made use of the Quantum Exact Simulation Toolkit (QuEST), and the recent development QuESTlink [40] which permits the user to use Mathematica as the integrated front end, and pyQuEST [41] which allows access to QuEST from Python. We are grateful to those who have contributed to all of these valuable tools. The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility [42] in carrying out this work and specifically the facilities made available from the EPSRC QCS Hub grant (agreement No. EP/T001062/1). The author also acknowledges funding from the EPSRC projects Robust and Reliable Quantum Computing (RoaRQ, EP/W032635/1) and Software Enabling Early Quantum Advantage (SEEQA, EP/Y004655/1).
@@ A.1 Quantum error mitigation
For this reason we consider applications where one assumes perfect unitary operations and the only source of error is due to, e.g., finite T count or discretisation of quantum rotation angles etc. Furthermore, operations $\mathcal{R}_{desired}$ and $\mathcal{U}_{l}$ are known exactly, hence arbitrary precision numerical representations can be used. This is all in contrast to QEM [27] whereby one approximately learns these from experiments and one thus needs to consider low-precision numerical representations.
@@ A.2 Probabilistic gate synthesis
Probabilistic gate sequences have been employed to improve the accuracy of Clifford+T sequences [43, 44, 45]. In particular, the desired unitary operation is approximated as $\mathcal{U}_{desired}\approx\sum_{l}p_{l}\mathcal{U}_{l}$ where $p_{l}$ is a probability distribution and $\mathcal{U}_{l}$ are is a set of operations close to $\mathcal{U}_{desired}$ but with finite T count. It was observed that the approximation error in $\sum_{l}p_{l}\mathcal{U}_{l}$ is always quadratically smaller than in $\mathcal{U}_{l}$. However, these techniques are approximate and as one increases the number of gates in the circuit, the fidelity of the quantum state decreases exponentially as demonstrated in [24].
@@ B.1 Proof of 1
$\mathbb{E}(\hat{\mathcal{U}}_{desired})=\sum_{l}p(l)\hat{\mathcal{U}}_{desired}=\sum_{l}p(l)\lVert\gamma\rVert_{1}\text{sign}(\gamma_{l})\mathcal{U}_{l},$
@@
$\mathbb{E}(\hat{\mathcal{U}}_{desired})=\sum_{l}\gamma_{l}\text{sign}(\gamma_{l})\mathcal{U}_{l}=\mathcal{U}_{desired}$
@@ B.2.1 Unbiasedness
B.2.2 Bounding the number of shots